# Beyond the Electric Dipole Approximation: Electric and Magnetic Multipole Contributions Reveal Biaxial Water Structure from SFG Spectra at the Air-Water Interface

### v2

@@ II.2 Spatially Resolved Second-Order Response Profile of the Bending Band
Figure 2 presents the position-dependent second-order response profile of the bending band, with $z$ given relative to the Gibbs dividing surface $z_{\mathrm{GDS}}$ [65]. Figure 2 (a) presents the mass density profile with a snapshot of the simulation box in the background. The density transitions from $1\text{\,}\mathrm{g}\mathrm{/}\mathrm{c}\mathrm{m}^{3}$ in bulk to zero over a range of about $5\text{\,}\mathrm{\SIUnitSymbolAngstrom}$, centered around $z_{\mathrm{GDS}}$.
@@
Figure 3: The orientation distribution function $\rho_{\mathrm{ORI}}(\theta,\psi)$ is shown for the same positions relative to $z_{\mathrm{GDS}}$ as in Figure 2 in (a)-(c). Three orientational species are distinguished: planar, pointing in, and pointing out, as illustrated at the top of panel (d). These orientations are marked with dots and contours in (a)-(c). Additionally, a snapshot of the simulation box viewed from the air is provided in (d), where molecules are color-coded according to their orientational species. (e) displays the fraction of each orientational species as a function of $z$. The molecular hyperpolarizability tensor $\tilde{\beta}_{yyz}(\theta,\psi)$, defined in Equation (3), is shown as a function of the molecular orientation in (f). The color indicates the intensity at bending-band frequencies, obtained by integrating the imaginary part $\tilde{\beta}^{\prime\prime}_{yyz}(\theta,\psi)$ using the same integration boundaries as in Equation (2). This intensity is decomposed into its uniaxial and biaxial components, defined in Equations (5) and (6) and shown in (g) and (h), respectively. The position-resolved profile of the bending band is decomposed into uniaxial and biaxial contributions in (i). The frequency-dependent coefficients $\tilde{\beta}^{lm}_{yyz}$ of $\tilde{\chi}^{(2,\mathrm{ORI})}_{yyz}$ defined in Equation (4)-(6) are presented in (j).
@@
We present the second-order electric dipole susceptibility $\tilde{\chi}^{(2,\mathrm{DL})}_{ijk}(z)$, which connects the average local electric E-field at the molecular centers to the resulting second-order electric dipole current density in Figure 2 (d) & (f). The superscript DL stands for dipole and local to distinguish it from regular susceptibilities, which are defined as the response of the macroscopic polarization density to macroscopic electric E-fields [66]. $\tilde{\chi}^{(2,\mathrm{DL})}_{ijk}(z)$ enables linking the macroscopic second-order electric dipole response to molecular orientation [67, 24, 25]. In defining $\tilde{\chi}^{(2,\mathrm{DL})}_{ijk}(z)$ via Equation (27), we account for the strength of the local E-field by dividing the second-order response of the electric dipole density $\tilde{s}^{(2,\mathrm{DD})}_{ijk}(z)$, by the local field factors, i.e. the ratios between the amplitudes of the local E-fields $\mathcal{E}^{\mathrm{L},\alpha}_{i}(z)$ acting on the molecular centers and the amplitudes of the external fields $\mathcal{F}^{\alpha}_{i}$, which correspond to D-fields for $i=z$ and to E-fields for $i=x$ or $y$. Here, $\alpha$ specifies the frequency of the three fields involved (SFG, VIS, IR). The relationship between $\tilde{s}^{(2,\mathrm{DD})}_{ijk}(z)$ and $\tilde{\chi}^{(2,\mathrm{DL})}_{ijk}(z)$ is given in Equation (28). The spatial integral over the second-order response profile $\tilde{s}^{(2,\mathrm{DD})}_{ijk}(z)$ in Equation (18) determines the contribution of the pure electric dipole to the experimentally measurable SFG spectrum $\tilde{S}^{(2,\mathrm{DD})}_{ijk}$ presented in Figure 1 (c). The local field factors are defined by
@@
The local field factor in the optical frequency range $f_{i}^{\mathrm{VIS}}(z)$ is presented in Figure 2 (b). Here, $f_{x}^{\mathrm{VIS}}(z)$ exhibits a line shape similar to the mass density profile. In bulk, we recover values dictated by the Lorentz field approximation [67, 25] $f^{\mathrm{VIS}}_{x}(-\infty)\approx(2+\varepsilon^{\mathrm{VIS}})/3$ and $f^{\mathrm{VIS}}_{z}(-\infty)\approx\left(2+\varepsilon^{\mathrm{VIS}}\right)/(3\varepsilon^{\mathrm{VIS}})$, as indicated by the horizontal dashed lines. Here, $\varepsilon^{\mathrm{VIS}}=1.77$ represents the optical dielectric constant of bulk water, extracted from the plateau value of the dielectric profile presented in Figure 5 (b). These findings are qualitatively consistent with the prediction of $f^{\mathrm{VIS}}_{i}(z)$ by Shiratori and Morita [68]. We compare the total second-order response function $\tilde{s}^{(2)}_{yyz}(z)$ defined in Equation (7) with the second-order electric dipole susceptibility $\tilde{\chi}^{(2,\mathrm{DL})}_{yyz}(z)$ defined in Equation (27), as well as its prediction based solely on molecular orientation, $\tilde{\chi}^{(2,\mathrm{ORI})}_{yyz}(z)$, in Figure 2 (e)-(g). It is evident that all profiles exhibit a positive contribution for values of $z-z_{\mathrm{GDS}}\geq 0$ and a negative contribution centered around $z-z_{\mathrm{GDS}}=$-2\text{\,}\mathrm{\SIUnitSymbolAngstrom}$\$. We calculate $\tilde{\chi}^{(2,\mathrm{ORI})}_{ijk}(z)$ by assigning the molecular hyperpolarizability tensor to each water molecule in its molecular frame, as defined in Equation (33), where the molecular hyperpolarizability tensor is extracted from a simulation of bulk water as described in SM Section VII [60]. By this, the only interface-specific input to $\tilde{\chi}^{(2,\mathrm{ORI})}_{ijk}(z)$ is the molecular orientation distribution, allowing us to test whether $\tilde{\chi}^{(2,\mathrm{DL})}_{ijk}(z)$ is a marker of the interfacial orientation anisotropy. To compare the amplitudes of $\tilde{\chi}^{(2,\mathrm{DL})}_{yyz}(z)$ and $\tilde{\chi}^{(2,\mathrm{ORI})}_{yyz}(z)$ we integrate over the bending band according to
$\bar{\chi}^{(2,\mathrm{DL})}_{ijk}(z)=c_{0}\int\limits_{\nu_{1}}^{\nu_{2}}\mathrm{d}\nu\,\tilde{\chi}^{{(2,\mathrm{DL}})\prime\prime}_{ijk}(z)$ (2)
@@
from $\nu_{1}=$1507\text{\,}\mathrm{c}\mathrm{m}^{-1}$\$ to $\nu_{2}=$1772\text{\,}\mathrm{c}\mathrm{m}^{-1}$\$. Here, $c_{0}$ is the speed of light in vacuum. We find quantitative agreement between $\bar{\chi}^{(2,\mathrm{DL})}_{yyz}(z)$ and $\bar{\chi}^{(2,\mathrm{ORI})}_{yyz}(z)$ in Figure 2 (c), demonstrating that $\bar{\chi}^{(2,\mathrm{DL})}_{yyz}(z)$ indeed primarily reflects interfacial orientation within the bending frequency range. The characteristic triple-layer structure of $\bar{\chi}^{(2,\mathrm{DL})}_{yyz}(z)$ will be discusses further below. By comparing $\tilde{\chi}^{(2,\mathrm{DL})}_{yyz}(z)$ (Fig. 2 (d) & (f)) with $\tilde{\chi}^{(2,\mathrm{ORI})}_{yyz}(z)$ (Fig. 2 (g)), it becomes clear that $\tilde{\chi}^{(2,\mathrm{ORI})}_{yyz}(z)$ does not reproduce the $z$-dependent frequency shift, present in $\tilde{\chi}^{(2,\mathrm{DL})}_{yyz}(z)$, which is thus due to the $z$-dependent anisotropic coordination of interfacial water.
@@ II.3 Spectral Signature of Biaxial Interfacial Water Ordering
After demonstrating that the second-order electric dipole susceptibility $\tilde{\chi}^{(2,\mathrm{DL})}_{yyz}$ serves as a quantitative probe of the interfacial orientation of molecules, we now analyze the interfacial orientation distribution. The orientation distribution function $\rho_{\mathrm{ORI}}(\theta,\psi)$, defined in Equation (34), is presented at three positions relative to $z_{\mathrm{GDS}}$ in Figure 3 (a)-(c). This function depends on the two Euler angles, $\theta$ and $\psi$. The angle $\theta$ describes the tilt of the molecular dipole axis relative to the surface normal: for $\theta=0$, the oxygen atom is oriented towards bulk water, and for $\theta=180^{\circ}$ towards air. The angle $\psi$ describes the molecular orientation around the dipole axis. The interfacial water structure is significantly biaxial, as revealed by the pronounced $\psi$-dependence of $\rho_{\mathrm{ORI}}(\theta,\psi)$. We distinguish three orientational species: the planar-oriented species (orange), where both OH-bond vectors lie in the interfacial plane, the inward-oriented species (magenta), where one OH-bond is orthogonal to the interfacial plane and pointing inward, and the outward-oriented species (cyan), where one OH-bond is orthogonal to the interfacial plane and pointing outward. These orientations are sketched on top of Figure 3 (d) and the corresponding Euler angles are marked with color-coded dots and contours (marking water molecules with angular deviation of $36^{\circ}$ from the idealized orientations) in the plots of $\rho_{\mathrm{ORI}}(\theta,\psi)$ in Figure 3 (a)-(c). A snapshot of the air-water interface viewed from air is presented in Figure 3 (d); here, the orientational species are color-coded. As observed in Figure 3 (a), at $2\text{\,}\mathrm{\SIUnitSymbolAngstrom}$ relative to $z_{\mathrm{GDS}}$, the inward-oriented species dominates; at $z=z_{\mathrm{GDS}}$, the planar species prevails (Fig. 3 (b)), and finally, close to the vapor region at $1.5\text{\,}\mathrm{\SIUnitSymbolAngstrom}$, the outward-oriented species becomes dominant (Fig. 3 (c)). We quantify this observation via the molecular number fraction profiles, which we present in Figure 3 (e). Here, we see that water is isotropic $5\text{\,}\mathrm{\SIUnitSymbolAngstrom}$ below $z_{\mathrm{GDS}}$, as the fraction of all orientational species is approximately $20\text{\,}\%$. The inward-oriented fraction peaks at $2\text{\,}\mathrm{\SIUnitSymbolAngstrom}$, at the location of the negative peak of $\tilde{\chi}^{(2,\mathrm{DL})}_{ijk}(z)$ in Figure 2 (f). The planar-oriented and outward-oriented species peak at $z_{\mathrm{GDS}}$ and near the vapor region at $2.5\text{\,}\mathrm{\SIUnitSymbolAngstrom}$, respectively. The uniaxial approximation neglects the biaxial water ordering, which corresponds to averaging the orientation distribution function over the angle $\psi$. To reveal the spectral effects of water biaxial ordering, we define the orientation and position dependent molecular hyperpolarizability $\tilde{\beta}_{ijk}(z,\theta,\psi)$ by
@@
where $\mu_{i}^{(2)}(t)$ is the second-order molecular electric dipole moment, $\mathcal{E}^{\mathrm{L,VIS}}_{j}(z)$ and $\mathcal{E}^{\mathrm{L,IR}}_{k}(z)$ are the amplitudes of the local E-fields acting on the molecular center at the position $z$, and $\theta$ and $\psi$ are the Euler angles, specifying the molecular orientation. Note that $\mu_{i}^{(2)}(t)$ is a source dipole that induces an additional linear response, as explained in SM Section IV B [60]. We extract $\tilde{\beta}_{yyz}(\theta,\psi)$, denoting the bulk water hyperpolarizability, from separate trajectories of bulk water, as described in SM Section VII [60]. In Figure 3 (f) we present the integral of the imaginary part $\tilde{\beta}^{\prime\prime}_{yyz}(\theta,\psi)$ over the same integration boundaries as in Equation (2), to quantify the orientation-dependent hyperpolarizability in the bending band. We decompose $\tilde{\beta}_{yyz}(\theta,\psi)$ into uniaxial and biaxial components according to
@@
shown in Figure 3 (g) & (h), respectively. Here, $q_{10}(\theta)=\cos{\theta}$, $q_{30}(\theta)=\frac{1}{2}\left[5\cos{{}^{3}\theta}-3\cos{\theta}\right]$ are the first and third Legendre polynomials specifying the dipole distribution and $q_{32}(\theta,\psi)=\cos{\theta}\sin{\theta}^{2}\left(\cos{{}^{2}\psi}-\sin{{}^{2}\psi}\right)$ accounts for the contribution due to the molecular biaxiality. Equations (4)–(6) express the rotation of the molecular hyperpolarizability into the laboratory frame in an exact way. The frequency-dependent imaginary parts of the coefficients $\tilde{\beta}^{10}_{yyz}$, $\tilde{\beta}^{30}_{yyz}$ and $\tilde{\beta}^{32}_{yyz}$ are presented in Figure 3 (j), where we observe that the biaxial contribution $\tilde{\beta}^{32}_{yyz}$ is significant. We observe that the negative area at $\theta>90^{\circ}$ in $\tilde{\beta}_{yyz}(\theta,\psi)$ in Figure 3 (f) does align well with the inward pointing orientation dominant at $2\text{\,}\mathrm{\SIUnitSymbolAngstrom}$ in Figure 3 (a) and the positive area at $\theta<90^{\circ}$ in Figure 3 (f) does align with the outwards pointing orientation dominant at $1.5\text{\,}\mathrm{\SIUnitSymbolAngstrom}$ in Figure 3 (c). We find that the uniaxial contribution in Figure 3 (g) for each orientational species cancels out, meaning that uniaxial water ordering does not lead by itself to an SFG signal in the bending region. In contrast, the biaxial component projects precisely out the difference between outwards- and inwards-oriented molecules (Fig. 3 (h)). We note that planar-oriented molecules are not detected by SFG spectroscopy because their mirror plane lies parallel to the interface, making them inversion symmetric. We dissect the integral over the bending band of $\tilde{\chi}^{(2,\mathrm{ORI})}_{yyz}(z)$, already presented in Figure 2 (c), into biaxial and uniaxial contributions in Figure 3 (i), demonstrating that $\tilde{\chi}^{(2,\mathrm{ORI})}_{yyz}(z)$ is dominated by the biaxiality of the orientation distribution.

### v3

@@ II.2 Spatially Resolved Second-Order Response Profile of the Bending Band
Figure 2 presents the depth-dependent second-order response profile of the bending band, relative to the Gibbs dividing surface position $z_{\mathrm{GDS}}$ [108]. Figure 2 (a) presents the mass density profile with a snapshot of the simulation box in the background. The density transitions from $1\text{\,}\mathrm{g}\mathrm{/}\mathrm{c}\mathrm{m}^{3}$ in bulk to essentially zero over a range of about $5\text{\,}\mathrm{\SIUnitSymbolAngstrom}$, centered around $z_{\mathrm{GDS}}$.
@@
Figure 3: The orientation distribution function $\rho_{\mathrm{ORI}}(\theta,\psi)$ is shown for the same positions relative to $z_{\mathrm{GDS}}$ as in Figure 2 in (a)-(c). Three orientational species are distinguished: planar, pointing in, and pointing out, as illustrated at the top of panel (d). These orientations are marked with dots and contours in (a)-(c). Additionally, a snapshot of the simulation box viewed from the air is provided in (d), where molecules are color-coded according to their orientational species. (e) displays the fraction of each orientational species as a function of $z$. The molecular hyperpolarizability tensor component $\tilde{\beta}_{yyz}(\theta,\psi)$, defined in Equation (4), is shown as a function of the molecular orientation in (f). The color indicates the intensity at bending-band frequencies, obtained by integrating the imaginary part $\tilde{\beta}^{\prime\prime}_{yyz}(\theta,\psi)$ using the same integration boundaries as in Equation (3). This intensity is decomposed into its uniaxial and biaxial components, defined in Equations (6) and (7) and shown in (g) and (h), respectively. The position-resolved profile of the bending band is decomposed into uniaxial and biaxial contributions in (i). The frequency-dependent coefficients $\tilde{\beta}^{lm\prime\prime}_{yyz}$ of $\tilde{\chi}^{(2,\mathrm{ORI})\prime\prime}_{yyz}$ defined in Equation (5)-(7) are presented in (j).
@@
We present the second-order electric dipole susceptibility $\tilde{\chi}^{(2,\mathrm{DL})}_{ijk}(z)$, which connects the average local electric E-field at the molecular centers to the resulting second-order electric dipole current density in Figure 2 (d) & (f). The superscript DL stands for dipole and local to distinguish it from regular susceptibilities, which are defined as the response of the macroscopic polarization density to macroscopic electric E-fields [22]. Though not directly experimentally measurable, $\tilde{\chi}^{(2,\mathrm{DL})}_{ijk}(z)$ is important since it links the macroscopic second-order electric dipole response to molecular orientation [6, 115, 88]. In defining $\tilde{\chi}^{(2,\mathrm{DL})}_{ijk}(z)$ via Equation (30), we account for the strength of the local E-field by dividing the second-order response of the electric dipole density $\tilde{s}^{(2,\mathrm{DD})}_{ijk}(z)$, by the local field factors, i.e. the ratios between the amplitudes of the local E-fields $\mathcal{E}^{\mathrm{L},\alpha}_{i}(z)$ acting on the molecular centers and the amplitudes of the external fields $\mathcal{F}^{\alpha}_{i}$, which correspond to D-fields for $i=z$ and to E-fields for $i=x$ or $y$. Here, $\alpha$ specifies the frequency of the three fields involved, i.e. $\alpha\in\{\mathrm{SFG},\mathrm{VIS},\mathrm{IR}\}$. The spatial integral over the second-order response profile $\tilde{s}^{(2,\mathrm{DD})}_{ijk}(z)$ in Equation (21) determines the pure electric dipole contribution to the experimentally measurable SFG spectrum $\tilde{S}^{(2,\mathrm{DD})}_{ijk}$; it is related to $\tilde{\chi}^{(2,\mathrm{DL})}_{ijk}(z)$ by
where $f^{\alpha}_{i}(z)$ are the laterally averaged local field factors, defined by
@@
The local field factors in the optical frequency range $f_{i}^{\mathrm{VIS}}(z)$ are presented in Figure 2 (b). Interestingly, $f_{x}^{\mathrm{VIS}}(z)$ exhibits a line shape similar to the mass density profile, while $f_{z}^{\mathrm{VIS}}(z)$ is markedly different. In bulk, we recover values dictated by the Lorentz field approximation [6, 88] $f^{\mathrm{VIS}}_{x}(-\infty)\approx(2+\tilde{\varepsilon}^{\mathrm{VIS}})/3$ and $f^{\mathrm{VIS}}_{z}(-\infty)\approx\left(2+\tilde{\varepsilon}^{\mathrm{VIS}}\right)/(3\tilde{\varepsilon}^{\mathrm{VIS}})$, as indicated by the horizontal dashed lines. Here, $\tilde{\varepsilon}^{\mathrm{VIS}}=1.77$ represents the optical dielectric constant of bulk water, extracted from the plateau value of the dielectric profile presented in Figure 5 (b). These findings are qualitatively consistent with the prediction of $f^{\mathrm{VIS}}_{i}(z)$ by Shiratori and Morita [116].
We compare the imaginary second-order response profile $\tilde{s}^{(2)\prime\prime}_{yyz}(z)$ defined in Equation (8) with the second-order electric dipole susceptibility $\tilde{\chi}^{(2,\mathrm{DL})\prime\prime}_{yyz}(z)$ defined in Equation (30), as well as its prediction based solely on molecular orientation, $\tilde{\chi}^{(2,\mathrm{ORI})\prime\prime}_{yyz}(z)$, in Figure 2 (e)-(g). It is evident that all profiles exhibit a positive contribution for values of $z-z_{\mathrm{GDS}}\geq 0$ (in red) and a negative contribution (in blue) centered around $z-z_{\mathrm{GDS}}=$-2\text{\,}\mathrm{\SIUnitSymbolAngstrom}$\$. We calculate $\tilde{\chi}^{(2,\mathrm{ORI})}_{ijk}(z)$ by summing the molecular hyperpolarizability tensor contributions assigned to each water molecule in its molecular frame, as defined in Equation (35), where the molecular hyperpolarizability tensor is extracted from a simulation of bulk water as described in SI Section VII. By this, the only interface-specific input to $\tilde{\chi}^{(2,\mathrm{ORI})}_{ijk}(z)$ is the molecular orientation distribution, allowing us to test whether $\tilde{\chi}^{(2,\mathrm{DL})}_{ijk}(z)$ is a marker of the interfacial orientation anisotropy. To compare the amplitudes of $\tilde{\chi}^{(2,\mathrm{DL})}_{yyz}(z)$ and $\tilde{\chi}^{(2,\mathrm{ORI})}_{yyz}(z)$ we integrate over the bending band according to
$\Sigma\tilde{\chi}^{(2,\mathrm{DL})\prime\prime}_{ijk}(z)=c_{0}\int\limits_{\nu_{1}}^{\nu_{2}}\mathrm{d}\nu\,\tilde{\chi}^{{(2,\mathrm{DL}})\prime\prime}_{ijk}(z,\nu)$ (3)
@@
from $\nu_{1}=$1507\text{\,}\mathrm{c}\mathrm{m}^{-1}$\$ to $\nu_{2}=$1772\text{\,}\mathrm{c}\mathrm{m}^{-1}$\$. Here, $c_{0}$ is the speed of light in vacuum. We find quantitative agreement between $\Sigma\tilde{\chi}^{(2,\mathrm{DL})\prime\prime}_{yyz}(z)$ and $\Sigma\tilde{\chi}^{(2,\mathrm{ORI})\prime\prime}_{yyz}(z)$ in Figure 2 (c), demonstrating that $\Sigma\tilde{\chi}^{(2,\mathrm{DL})\prime\prime}_{yyz}(z)$ indeed primarily reflects interfacial orientation within the bending frequency range. The characteristic triple-layer structure of $\Sigma\tilde{\chi}^{(2,\mathrm{DL})\prime\prime}_{yyz}(z)$ will be discusses further below. By comparing $\tilde{\chi}^{(2,\mathrm{DL})\prime\prime}_{yyz}(z)$ (Fig. 2 (d) & (f)) with $\tilde{\chi}^{(2,\mathrm{ORI})\prime\prime}_{yyz}(z)$ (Fig. 2 (g)), it becomes clear that $\tilde{\chi}^{(2,\mathrm{ORI})\prime\prime}_{yyz}(z)$ does not reproduce the $z$-dependent frequency shift, present in $\tilde{\chi}^{(2,\mathrm{DL})\prime\prime}_{yyz}(z)$, which is thus due to the $z$-dependent anisotropic coordination of interfacial water.
@@ II.3 Spectral Signature of Biaxial Interfacial Water Ordering
After demonstrating that the second-order electric dipole susceptibility $\tilde{\chi}^{(2,\mathrm{DL})}_{yyz}$ serves as a quantitative probe of the interfacial orientation of molecules, we now analyze the interfacial orientation distribution.
The orientation distribution function $\rho_{\mathrm{ORI}}(\theta,\psi)$, defined in Equation (36), is presented at three depth $z-z_{\mathrm{GDS}}$ in Figure 3 (a)-(c). This function depends on the two Euler angles, $\theta$ and $\psi$. The angle $\theta$ describes the tilt of the molecular dipole axis relative to the surface normal: for $\theta=0$, the oxygen atom is oriented towards bulk water, and for $\theta=180^{\circ}$ towards air. The angle $\psi$ describes the molecular orientation around the dipole axis. The interfacial water structure is significantly biaxial, as revealed by the pronounced $\psi$-dependence of $\rho_{\mathrm{ORI}}(\theta,\psi)$.
We distinguish three orientational species: the planar-oriented species (orange), where both OH-bond vectors lie in the interfacial plane, the inward-oriented species (magenta), where one OH-bond is orthogonal to the interfacial plane and pointing inward, and the outward-oriented species (cyan), where one OH-bond is orthogonal to the interfacial plane and pointing outward. These orientations are sketched on top of Figure 3 (d) and the corresponding Euler angles are marked with color-coded dots and contours (marking water molecules with angular deviation of $36^{\circ}$ from the idealized orientations) in the plots of $\rho_{\mathrm{ORI}}(\theta,\psi)$ in Figure 3 (a)-(c). A snapshot of the air-water interface viewed from air is presented in Figure 3 (d); here, the orientational species are color-coded.
As observed in Figure 3 (a), at $z-z_{\mathrm{GDS}}=$-2\text{\,}\mathrm{\SIUnitSymbolAngstrom}$\$, the inward-oriented species dominates; at $z=z_{\mathrm{GDS}}$, the planar species prevails (Fig. 3 (b)), and finally, close to the vapor region at $z-z_{\mathrm{GDS}}=$1.5\text{\,}\mathrm{\SIUnitSymbolAngstrom}$\$, the outward-oriented species becomes dominant (Fig. 3 (c)). We quantify this observation via the molecular number fraction profiles, which we present in Figure 3 (e). Here, we see that water is isotropic $5\text{\,}\mathrm{\SIUnitSymbolAngstrom}$ below $z_{\mathrm{GDS}}$, as the fraction of all orientational species is approximately $20\text{\,}\%$. The inward-oriented fraction peaks at $z-z_{\mathrm{GDS}}=$-2\text{\,}\mathrm{\SIUnitSymbolAngstrom}$\$, at the location of the negative peak of $\tilde{\chi}^{(2,\mathrm{DL})}_{ijk}(z)$ in Figure 2 (f). The planar-oriented and outward-oriented species peak at $z=z_{\mathrm{GDS}}$ and near the vapor region at $z-z_{\mathrm{GDS}}=$2.5\text{\,}\mathrm{\SIUnitSymbolAngstrom}$\$, respectively.
The uniaxial approximation, which corresponds to averaging the orientation distribution function over the angle $\psi$, neglects the biaxial water ordering. To reveal the spectral effects of water biaxial ordering, we define the orientation and position dependent molecular hyperpolarizability $\tilde{\beta}_{ijk}(z,\theta,\psi)$ by
@@
where $\mu_{i}^{(2)}(t)$ is the second-order molecular electric dipole moment, $\mathcal{E}^{\mathrm{L,VIS}}_{j}(z)$ and $\mathcal{E}^{\mathrm{L,IR}}_{k}(z)$ are the amplitudes of the local E-fields acting on the molecular center at the position $z$, and $\theta$ and $\psi$ are the Euler angles, specifying the molecular orientation. Note that $\mu_{i}^{(2)}(t)$ is a source dipole that induces an additional linear response, as explained in SI Section IV B. We extract $\tilde{\beta}_{yyz}(\theta,\psi)$, from simulation trajectories of bulk water, as described in SI Section VII. In Figure 3 (f) we present the integral of the imaginary part $\tilde{\beta}^{\prime\prime}_{yyz}(\theta,\psi)$ over the same integration boundaries as in Equation (3), to quantify the orientation-dependent hyperpolarizability in the bending band. We decompose $\tilde{\beta}_{yyz}(\theta,\psi)$ into uniaxial and biaxial components according to
@@
shown in Figure 3 (g) & (h), respectively. Here, $q_{10}(\theta)=\cos{\theta}$, $q_{30}(\theta)=\frac{1}{2}\left[5\cos{{}^{3}\theta}-3\cos{\theta}\right]$ are the first and third Legendre polynomials specifying the dipole distribution and $q_{32}(\theta,\psi)=\cos{\theta}\sin{\theta}^{2}\left(\cos{{}^{2}\psi}-\sin{{}^{2}\psi}\right)$ accounts for the contribution due to the molecular biaxiality. Equations (5)–(7) express the rotation of the molecular hyperpolarizability into the laboratory frame in an exact way. The frequency-dependent imaginary parts of the coefficients $\tilde{\beta}^{10}_{yyz}$, $\tilde{\beta}^{30}_{yyz}$ and $\tilde{\beta}^{32}_{yyz}$ are presented in Figure 3 (j), where we observe that the biaxial contribution $\tilde{\beta}^{32}_{yyz}$ is significant.
We observe that the negative area (blue) at $\theta>90^{\circ}$ in $\tilde{\beta}_{yyz}(\theta,\psi)$ in Figure 3 (f) aligns with the inward pointing orientation dominant at $2\text{\,}\mathrm{\SIUnitSymbolAngstrom}$ in Figure 3 (a) and the positive area (red) at $\theta<90^{\circ}$ in Figure 3 (f) aligns with the outwards pointing orientation dominant at $1.5\text{\,}\mathrm{\SIUnitSymbolAngstrom}$ in Figure 3 (c). As seen in Figure 3 (g), the uniaxial contribution changes sign within the domain of each orientational species, suggesting that uniaxial water ordering by itself does not lead to a SFG signal in the bending region, as will be rigorously demonstrated below. In contrast, the biaxial component projects precisely out the difference between outward- and inward-oriented molecules (Fig. 3 (h)). We note that planar-oriented molecules are not detected by SFG spectroscopy (neither uniaxial nor biaxial), because their mirror plane lies parallel to the interface, making them inversion symmetric. We dissect the integral over the bending band $\Sigma\tilde{\chi}^{(2,\mathrm{ORI})\prime\prime}_{yyz}(z)$, already presented in Figure 2 (c), into biaxial and uniaxial contributions in Figure 3 (i), demonstrating that $\Sigma\tilde{\chi}^{(2,\mathrm{ORI})\prime\prime}_{yyz}(z)$ is dominated by the biaxiality of the orientation distribution. Thus, our analysis of the simulation trajectories shows that interfacial water is significantly biaxial and that the sign of the bending-mode dipole response is an unambiguous indicator of the relative populations of biaxially ordered inward- and outward-oriented water molecules illustrated in Figure 3 (d).